# A New Causal Rule Learning Approach to Interpretable Estimation of Heterogeneous Treatment Effect

### v2

@@ 3.2.1 Rule discovery via causal forest
Our model utilizes causal forests [45] to derive the initial set of potential rules and corresponding CATE estimates. Causal forest is a non-parametric method for heterogeneous treatment effect estimation whose estimator has been shown to have good statistical properties like point-wise consistency and asymptotical normality for the true CATE under the three Neyman-Rubin assumptions [45].
A causal forest consists of many causal trees [3], each trained using a random subsample of observations and covariates. Given a causal tree, it is easy and intuitive to decompose the tree into multiple rules: the path from the root to a leaf node naturally forms a rule formulated in definition (1). Instances within a leaf node directly make up a potential subpopulation. Further, we use the following plug-in estimator of difference-in-means as an estimate for the corresponding CATE of a given subgroup $G$ in a leaf node:
@@ 3.2.2 Rule selection with the D-learning method
Due to the randomization and greediness introduced by forest-based methods, fake and redundant rules are inevitably generated [46]. The rule selection process aims to filter out those irrelevant and unnecessary rules from the initial rule set to improve the accuracy of ITE estimation and size down the rule set for interpretability purposes.
@@
For this end, we leverage the D-learning method [34] to learn a sparse combination of rules that can estimate ITE accurately. The original D-learning method aims to directly learn the optimized individual treatment rule (ITR) $d^{*}(X)$ in a single step without model specification, defined as:
$d^{*}(X)=\textnormal{sign}\Big{(}\mathbf{E}\Big{[}\frac{YA}{\pi(A,X)}|X\Big{]}\Big{)}=\textnormal{sign}(\tau^{*}(X)),$ (4)
@@
where the expectation $\mathbf{E}[YA/\pi(A,X)]$ or $\tau^{*}(X)$ is the exact desired ITE. (4) leverages the propensity score defined as the probability of receiving treatment $a$ given $X$, i.e., $\pi(A,X)=Pr(A|X)$. For RCT data, we assume this score is already known. For observational data, we have to specify a proper model and learn the score from the data.
@@
It has been proved in [34] that under the assumption of interchange between differentiation and expectation,
$\tau^{*}(X)\in\mathop{\arg\min}_{\tau(X)}\mathbf{E}\Big{[}\frac{(2YA-\tau(X))^{2}}{\pi(A,X)}|X\Big{]},$
which transforms the estimation of unobserved ITE into an expectation minimization problem where $\tau(X)$ can be any proper form, allowing a flexible specification no matter linear or nonlinear.
Therefore, replacing $\tau(X)$ with a linear combination of subgroup CATE, i.e., the right-hand side of (2) and imposing additional $\ell_{1}$ penalty on the weights of the rules to ensure sparsity, we can reformulate the estimation of $\tau^{*}(X)$ as a $\ell_{1}$ penalized regression problem and obtain the least absolute shrinkage and selection operator (LASSO) estimator [41] of the rule weights $\hat{\mathbf{\beta}}^{lasso}$:
$\hat{\mathbf{\beta}}^{lasso}=\arg\min_{\mathbf{\beta}}\Big{\{}\frac{1}{2N}\sum^{N}_{i=1}\frac{1}{\pi(A_{i},X_{i})}\{2Y_{i}A_{i}-\sum_{m=1}^{M}\beta_{m}\hat{\tau}_{m}\hat{r}_{m}(X_{i})\}^{2}+\lambda\sum_{m=1}^{M}|\beta_{m}|\Big{\}},$ (5)
where $\lambda>0$ is the tuning parameter that controls the degree of penalty. Then we can easily obtain CRL estimates of ITE: $\hat{\tau}_{CRL}(X)=\sum_{m=1}^{M}\hat{\beta}^{lasso}_{m}\hat{\tau}_{m}\hat{r}_{m}(X)$.
@@
Just as [31] states, when we impose sparsity on our model, we limit the number of non-zero parameters in the model and interpret the rules corresponding to those parameters (based on sign and magnitude) as meaningful to the outcome of interest. Since a large number of weights are zero (i.e., $\hat{\beta}^{lasso}$ usually has many elements equal to 0), their corresponding rules are actually removed from the candidate set and make no contribution to our estimator. Imposing this sparsity of rules also highlights the distinction between our CRL estimator and the causal forest estimator ,which simply takes the average over all the tree estimates (subgroup CATEs in our method) as the final estimates, while our method learns a weighted combination of those estimates. In other words, we reweight the tree estimates instead of using them equally. In this sense, causal forest estimates can be viewed as a special case of our method. We show in Figure 4(d) that reweighting the rules with the D-learning method is able to eliminate a large proportion of unwanted rules causal forest generates.
@@ 3.2.3 Rule analysis
Overall analysis. Suppose D-learning selects a rule set that consists of $n_{1}$ rules ($n_{1}\leq M$) whose weights are non-zero in model (5), i.e., $\hat{R}_{1}=\{\hat{r}_{1},\hat{r}_{2},\dots,\hat{r}_{n_{1}}\}$ (for simplicity, we use $\hat{r}_{j}$ instead of $\hat{r}_{j}(x)$ to represent a rule). The corresponding CRL estimates of ITE $\hat{\tau}_{CRL}$ are also obtained based on this rule set. As estimation accuracy is the basis of interpretability, the overall analysis aims to evaluate how this rule set collectively performs on ITE estimation using various performance metrics. For simulated data where the true treatment effect is already known, we define Mean Squared Error (MSE), Mean Potential Outcomes (MPO) and Population Overlap (PO) to evaluate and compare our estimator with other baselines. For observational data where the ground truth can never be observed, the above metrics can no longer be used, so we use ITE-based prediction accuracy, Empirical Expected Outcome (EEO) and Treatment Efficient Frontier. Details of these metrics can be found in section 4.1 for simulated data and 4.2.2 for real-world data.
@@
Suppose the above overall analysis shows that our method is able to give comparable ITE estimates to that of baseline methods, we then continue the following analysis, which evaluates each rule in terms of i) the ability to distinguish different levels of ITE, i.e., rule significance, and ii) the role of each component in the rule, i.e., rule decomposition analysis.
@@
Rule significance analysis. A rule defines two subpopulations: the population that satisfies the rule $G^{Y}=\{i|r(X_{i})=1\}$ and the rest that does not $G^{N}=\{i|r(X_{i})=0\}$. Therefore, we want to test the discriminating power of a rule by checking if $G^{Y}$ and $G^{N}$ differ significantly in the magnitude of ITE estimates, using the well-known two-sample Kolmogorov–Smirnov test [7]. This test is commonly used to determine whether two samples originate from the same continuous distribution. In our context, the null hypothesis posits that the ITE estimates for groups $G^{Y}$ and $G^{N}$ derive from an identical distribution. If the null hypothesis cannot be rejected, we assume there is no significant difference in the ITE estimates between the two groups. Specifically, a $p$-value $\leq 0.05$ implies that the rule can identify a subgroup with significantly different treatment effects than the rest of the population. We perform this test on each rule in $\hat{R}_{1}$ and remove those with $p$-value $>0.05$. The remaining $n_{2}$ rules ($n_{2}\leq n_{1}$) form the new candidate rule set $\hat{R}_{2}$.
Rule decomposition analysis. Decomposition analysis evaluates how each component in a rule contributes to the complete rule in differentiating treatment effect. For example, if we have a rule $\hat{r_{j}}$ that has three components $a$, $b$ and $c$, we can derive three revised rules: $\hat{r}^{a}_{j}$, $\hat{r}^{b}_{j}$ and $\hat{r}^{c}_{j}$. $\hat{r}^{a}_{j}$ is the rule with component $a$ removed from the original rule, that is, the rule formed only by component $b$ and $c$ and similarly for $\hat{r}^{b}_{j}$ and $\hat{r}^{c}_{j}$. We do the above significance analysis on all revised rules and obtain their corresponding $p$-values $p^{a}_{j}$, $p^{b}_{j}$ and $p^{c}_{j}$. We think a higher $p$-value of the revised rule than that of the original complete rule, say, $p^{c}_{j}>p_{j}$, implies that the revised rule $\hat{r}^{c}_{j}$ is less significant in distinguishing subpopulations with different levels of ITE estimates than the original entire rule $\hat{r_{j}}$. Hence, the removed component $c$ contributes to the significance of the ITE difference. The higher the $p$-value of the revised rule, the more critical the variable being removed is, which reveals the variable most accountable for the group’s heterogeneity. In contrast, a lower value suggests the corresponding component(s) may be unnecessary and meaningless to the complete rule since its removal does not reduce the significance of the difference but increases it. Therefore, if all revised rules have higher $p$-values ($p_{j}<$ min$\{p^{a}_{j},p^{b}_{j},p^{c}_{j}\}$), then all the components in the rule work synergistically to contribute to its discriminating power. In this step, we retain the rules that have higher $p$-values on all the corresponding revised rules to form a new refined set $\hat{R}_{3}$ with $n_{3}$ rules ($n_{3}\leq n_{2}$).
@@
The rule analysis step shows a universal way to analyze the identified rules through different levels. However, these tools highly rely on the data in hand. Therefore, we highly recommend using this analysis procedure as a screening step for the lengthy list of rules and picking out the top subgroups for more rigorous statistical test and further interpretation.
@@
Our whole workflow helps us to estimate and build a better understanding of HTE for complex diseases by finding potential subgroups of heterogeneity and exploring the relationship between ITE and CATEs of these subgroups. Take Figure 1 as an example, the rule discovery step first identifies a large pool of interpretable rules as candidate subgroups. The rule selection process then removes a great proportion of non-informative and fake subgroups from the candidate set and uses the remaining $M$ subgroups as the selected sources of effect heterogeneity to learn the connection between ITE and subgroup CATEs (through $\tau$ and $\beta$). Finally, the rule analysis part evaluates how accurate these $M$ subgroups estimate ITE as a whole and looks deeper into each of the $M$ subgroups to evaluate how each subgroup and the factors involved in the subgroup contributes to HTE estimation.

### v3

@@ 3.2.1 Rule discovery via causal forest
Our model utilizes causal forests [wager2018estimation] to derive the initial set of potential rules and corresponding CATE estimates. Causal forest is a non-parametric method for heterogeneous treatment effect estimation whose estimator has been shown to have good statistical properties like point-wise consistency and asymptotic normality for the true CATE under the three Neyman-Rubin assumptions [wager2018estimation].
A causal forest consists of many causal trees [athey2016recursive], each trained using a random subsample of observations and covariates. Given a causal tree, it is easy and intuitive to decompose the tree into multiple rules: the path from the root to a leaf node naturally forms a rule formulated in definition (1). Instances within a leaf node directly make up a potential subpopulation. Further, we use the following plug-in estimator of difference-in-means as an estimate for the corresponding CATE of a given subgroup $G$ in a leaf node:
@@ 3.2.2 Rule selection with the D-learning method
Given the $\{\hat{r}_{m},\hat{\tau}_{m}\}^{M}_{m=1}$ obtained from last step and our linear assumption on the $\tau(X)$ in (2), we provide our estimation of ITE formulated as
$\hat{\tau}(X)=\sum_{m=1}^{M}\beta_{m}\hat{\tau}_{m}\hat{r}_{m}=\left(\hat{\tau}\circ\hat{r}(X)\right)^{\top}\beta:=\tilde{X}^{\top}\beta,$ (4)
where $\circ$ is the Hadamard product of $\hat{r}(X)=\left(\hat{r}_{1}(X),\cdots,\hat{r}_{M}(X)\right)^{\top}\in\mathbb{R}^{M}$ and $\hat{\tau}=\left(\hat{\tau}_{1},\cdots,\hat{\tau}_{M}\right)^{\top}\in\mathbb{R}^{M}$. $\beta=(\beta_{1},\beta_{2},\cdots,\beta_{M})^{\top}$. This is merely a transformation from the original sample vector $X$ to the modified rule space $\tilde{X}\in\mathbb{R}^{M}$.
@@
Due to the randomization and greediness introduced by forest-based methods, fake and redundant rules are inevitably generated [wang2020an]. Therefore, the rule selection process aims to filter out these irrelevant and unnecessary rules from the initial rule set through the learning of the coefficients $\beta$ to improve the accuracy of ITE estimation and size down the rule set for interpretability purposes. For this end, we leverage the D-learning method [qi2018d] to learn a sparse combination of rules that can estimate ITE accurately.
@@
The original D-learning method aims to directly learn the optimized individual treatment decision $d^{*}(X)$ in a single step without model specification, defined as:
$d^{*}(X)=\textnormal{sign}\Big(\mathbf{E}\Big[\frac{YA}{\pi(A,X)}|X\Big]\Big)=\textnormal{sign}(f_{0}(X)),$ (5)
@@
where the $f_{0}(X):=\mathbf{E}\Big[\frac{YA}{\pi(A,X)}|X\Big]$ is the desired true $\tau(X)$. (5) leverages the propensity score defined as the probability of receiving treatment $a$ given $X$, i.e., $\pi(A,X)=Pr(A|X)$. For RCT data, we assume this score is already known. For observational data, we have to specify a proper model and learn the score from the data.
@@
It has been demonstrated in [qi2018d] that under the assumption of interchange between differentiation and expectation,
$f_{0}(X)\in\mathop{\arg\min}_{f}\mathbf{E}\Big[\frac{(2YA-f(X))^{2}}{\pi(A,X)}|X\Big],$ (6)
which transforms the estimation of unobserved true ITE into an expectation maximization problem where $f(X)$ can be any proper form, allowing a flexible specification, whether linear or nonlinear. Therefore, we replace $f(X)$ with (4) and impose additional $\ell_{1}$ penalty on $\beta$ to ensure sparsity. Then the estimation of $f_{0}(X)$ is reformulated as an $\ell_{1}$ penalized regression problem, and we can obtain the least absolute shrinkage and selection operator (LASSO) estimator [tibshirani1996regression] of $\beta$:
$\hat{\mathbf{\beta}}=\arg\min_{\mathbf{\beta}}\Big\{\frac{1}{2N}\sum^{N}_{i=1}\frac{1}{\pi(A_{i},X_{i})}\{2Y_{i}A_{i}-\tilde{X}_{i}^{\top}\beta\}^{2}+\lambda\|\beta\|_{1}\},$ (7)
where $\lambda\geq 0$ is the tuning parameter.
@@
Just as [yubinPDR2019] states, when we impose sparsity on our model, we limit the number of non-zero parameters in the model and interpret the rules corresponding to those parameters (based on sign and magnitude) as meaningful to the outcome of interest. Since a large number of weights are zero (i.e., $\hat{\beta}$ usually has many elements equal to 0), their corresponding rules are actually removed from the candidate set and do not contribute to our estimator. Imposing this sparsity of rules also highlights the distinction between our CRL estimator and the causal forest estimator, which simply takes the average over all the tree estimates (subgroup CATEs in our method) as the final estimates, while our method learns a weighted combination of those estimates. In other words, we reweight the tree estimates instead of using them equally. In this sense, causal forest estimates can be viewed as a special case of our method. We show in Figure 4(d) that reweighting the rules with the D-learning method can eliminate a large proportion of unwanted rules that the causal forest generates.
[…]
$V(d)$ $=\mathbf{E}[Y|A=d(\mathbf{X})],$
$V(\hat{d}_{N})$ $=\mathbf{E}[Y|A=\hat{d}_{N}(\mathbf{X})]$
where $\mathbf{X}\in\mathbb{R}^{N\times p}$ is the original design matrix composed of $N$ samples and $p$ covariates. Both $V(d)$ and $V(\hat{d}_{N})$ yield an $N$-dimensional vector for every sample.
With the above definitions, we show below that convergence rates of the expected outcome under our treatment decision, i.e., $V(\hat{d}_{N})$, can be achieved when the following assumptions and conditions hold:
  1. 1.
    The compatibility condition. For some constant $\phi(S)$ and any vector $\beta\in\mathbb{R}^{M}$ with $\|\beta_{S^{c}}\|_{1}\leq 3\|\beta_{S}\|_{1}$, we have
    $\|\beta_{S}\|_{1}^{2}\leq\left(\beta^{\top}\hat{\Sigma}\beta\right)|S|/\phi^{2}(S),$ (8)
    where $\beta_{S}$ denotes the subvector of $\beta$ consisting of the components indexed by the set $S$, $\beta_{S^{c}}$ is the complement of $\beta_{S}$, $\hat{\Sigma}=\frac{1}{N}\tilde{\mathbf{X}}_{N}^{\top}\tilde{\mathbf{X}}_{N}$, and $\tilde{\mathbf{X}}_{N}=\left(\tilde{X}_{1},\cdots,\tilde{X}_{N}\right)^{\top}\in\mathbb{R}^{N\times M}$. When the parameter is the oracle $\beta^{*}$, we use $s_{*}$ and $\phi_{*}$ to replace $|S_{*}|$ and $\phi(S_{*})$ where $S_{*}=\{j:\beta_{j}^{*}\neq 0\}$ for simplicity.
  2. 2.
    Assume $Y_{i}=m(X_{i})+A_{i}\delta(X_{i})+W_{i}$, where $m(X_{i})$ is the main effect of covariates $X_{i}$ for both treatments and $W_{i}$ is the mean zero random error term with its variance $\sigma^{2}>0$. Let $\epsilon_{i}=2R_{i}A_{i}-\mathbf{E}\left[\frac{R_{i}A_{i}}{\pi(A_{i},X_{i})}\mid X_{i}\right]$ for $i=1,2,\cdots,N$.
  3. 3.
    Assume $m(X_{i})=X_{i}^{\top}\gamma_{0}$, where $\|\gamma_{0}\|\leq\mathcal{O}\left(\sqrt{\log(2M)}\right)$.
  4. 4.
    $\max_{i\in[M]}|\tau_{i}|\leq a$ for some $a>0$.
  5. 5.
    There exists some $\rho>0$ such that $\gamma^{\top}{\Sigma}\gamma\leq\rho\|\gamma\|_{2}^{2}$ for any $\gamma$, where $\Sigma=\frac{1}{N}\mathbf{X}_{N}^{\top}\mathbf{X}_{N}$ and $\mathbf{X}_{N}$ is the original design matrix.
  6. 6.
    Define the oracle $\mathbf{\beta}^{*}$ as $\beta^{*}=\mathrm{argmin}_{\beta}\{\|\tilde{\mathbf{X}}_{N}\beta-f_{0}({\mathbf{X}}_{N})\|_{2}^{2}/N+\frac{4\lambda^{2}s_{\boldsymbol{\beta}}}{\phi^{2}(S_{\boldsymbol{\beta}})}\}$, then assume $\|\tilde{\mathbf{X}}_{N}\boldsymbol{\beta}^{*}-f_{0}({\mathbf{X}}_{N})\|_{2}^{2}/N\leq\lambda\|\hat{\boldsymbol{\beta}}_{S_{*}}-\boldsymbol{\beta}_{S_{*}}^{*}\|_{1}$ where $S_{*}=\{j:\beta_{j}^{*}\neq 0\}$, $\hat{\beta}_{j,S_{*}}=\hat{\beta}_{j}\mathbb{I}(j\in S_{*})$.
  7. 7.
    The general Margin condition (gMC). For any $\epsilon>0$ there exists some gMC constant $C>0$ and $\alpha>0$, such that $\mathbb{P}(|f_{0}(X)|\leq\epsilon)\leq C\epsilon^{\alpha}$.
Specifically, Assumption (1) is a standard condition in the LASSO literature, ensuring that the $\ell_{1}$-norm of the coefficients on the true support can be controlled by the prediction norm. This property is crucial for establishing oracle inequalities and consistency results (see [vandeGeer2009oracle]). Assumptions (2)–(5) pertain to the modeling of the reward function $Y_{i}$ and impose restrictions on the original sample space, following [qi2018d]. Assumption (6) is a technical condition introducing an oracle linear parameter $\beta^{*}$ and describing its approximation behavior to the true decision function $f_{0}$, as also considered in [qi2018d]. Finally, Assumption (7) characterizes the distributional behavior of the decision function near zero; larger values of $\alpha$ correspond to a larger exponent $1+\alpha/(2+\alpha)$, which in turn yields a sharper upper bound in (10).
Then, we provide the following main theorem:
Theorem 3.1.
Let the tuning parameter $\lambda$ be
$\lambda=16\sqrt{2}t^{2}\sqrt{\frac{\log^{2}(2M)}{N}}.$ (9)
If the true treatment decision function is linear, that is, $f_{0}=\tilde{\mathbf{X}}_{N}{\beta}^{*}$, then when the assumptions (1)-(7) hold, with probability at least $1-\frac{C}{t^{2}}$ where $C$ depends on $a,\rho,\sigma$, we have
$\|V(d^{*})-V(\hat{d}_{N})\|_{2}\leq C_{2}\left(\frac{\log(2M)}{N}\right)^{\frac{1+\alpha}{2+\alpha}},$ (10)
where $C_{2}$ is determined by the gMC constant, $t$, $\phi_{S_{*}}$, and $|S_{*}|$.
Remark 1.
This theorem provides the explicit convergence rate of the value difference to zero when the true ITE function is indeed linear in the transformed sample space, i.e., $f_{0}=\tilde{\mathbf{X}}_{N}\beta^{*}$. The rate $(\log(2M)/N)^{\frac{1+\alpha}{2+\alpha}}$ depends on the number of candidate rules $M$ only through its logarithm, justifying the use of a large initial rule pool generated by the causal forest. The exponent $\frac{1+\alpha}{2+\alpha}$, governed by the gMC constant $\alpha$, highlights the beneficial effect of a margin condition—stronger separation between treatment effects (larger $\alpha$) leads to faster convergence of the policy value. This theoretical guarantee reinforces the practical utility of the CRL framework. The proof of Theorem 3.1 can be found in Appendix C.
@@ 3.2.3 Rule analysis
Overall analysis. Suppose D-learning selects a rule set that consists of $n_{1}$ rules ($n_{1}\ll M$) whose weights are non-zero in model (7), i.e., $\hat{R}_{1}=\{\hat{r}_{1},\hat{r}_{2},\dots,\hat{r}_{n_{1}}\}$ (for simplicity, we use $\hat{r}_{j}$ instead of $\hat{r}_{j}(x)$ to represent a rule). The corresponding CRL estimates of ITE $\hat{\tau}_{CRL}$ are also obtained based on this rule set. As estimation accuracy is the basis of interpretability, the overall analysis aims to evaluate how this rule set collectively performs on ITE estimation using various performance metrics. For simulated data where the true treatment effect is already known, we define Mean Squared Error (MSE), Mean Potential Outcomes (MPO), and Population Overlap (PO) to evaluate and compare our estimator with other baselines. For observational data where the ground truth can never be observed, the above metrics can no longer be used, so we use ITE-based prediction accuracy, Empirical Expected Outcome (EEO), Mean-squared Prediction Error (MSPE) of transformed outcome, and Treatment Efficient Frontier. Details of these metrics can be found in section 4.1 for simulated data and 4.4.2 for real-world data.
@@
Suppose the above overall analysis shows that our method is able to give comparable ITE estimates to those of baseline methods, we then continue the following analysis, which evaluates each rule in terms of i) the ability to distinguish different levels of ITE, i.e., rule significance, and ii) the role of each component in the rule, i.e., rule decomposition analysis.
@@
Rule significance analysis. A rule defines two subpopulations: the population that satisfies the rule $G^{Y}=\{i|r(X_{i})=1\}$ and the rest that does not $G^{N}=\{i|r(X_{i})=0\}$. Therefore, we want to test the discriminating power of a rule by checking if $G^{Y}$ and $G^{N}$ differ significantly in the magnitude of ITE estimates, using the well-known two-sample Kolmogorov–Smirnov test [berger2014kolmogorov]. This test is commonly used to determine whether two samples originate from the same continuous distribution. In our context, the null hypothesis posits that the ITE estimates for groups $G^{Y}$ and $G^{N}$ derive from an identical distribution. If the null hypothesis cannot be rejected, we assume there is no significant difference in the ITE estimates between the two groups. Specifically, for a given significance level $\alpha$, a $p$-value $\leq\alpha$ implies that the rule can identify a subgroup with significantly different treatment effects than the rest of the population. We perform this test on each rule in $\hat{R}_{1}$ and remove those with $p$-value $\alpha$. To avoid a high false discovery rate (FDR) in this multiple testing, we recommend using FDR control techniques such as the well-known Benjamini-Hochberg method [benjamini_controlling_2018] and removing rules based on these FDR-adjusted $p$-values instead of the original ones. The choice of $\alpha$ and the FDR level $q$ is up to domain experts and should be justified rigorously. In our experiments below, we simply adopt the conventional choice of 0.05 for both thresholds for illustrative purposes. The remaining $n_{2}$ rules ($n_{2}\leq n_{1}$) form the new candidate rule set $\hat{R}_{2}$.
[…]
@@
In conclusion, the rule analysis step shows a universal way to analyze the identified rules through different levels. However, these tools highly rely on the data at hand. Therefore, we highly recommend using this analysis procedure as a screening step for the lengthy list of rules and picking out the top subgroups for more rigorous statistical tests and further interpretation.
3.3 Discussion on the workflow design of CRL
@@
As presented in the last section, our whole workflow helps us to estimate and build a better understanding of HTE for complex diseases by finding potential subgroups of heterogeneity and exploring the relationship between ITE and CATEs of these subgroups. As Figure 2 shows, the rule discovery step first identifies a large pool of interpretable rules as candidate subgroups. The rule selection process then removes a great proportion of non-informative and fake subgroups from the candidate set and uses the remaining $M$ subgroups as the selected sources of effect heterogeneity to learn the connection between ITE and subgroup CATEs (through $\tau$ and $\beta$). Finally, the rule analysis part evaluates how accurate these $M$ subgroups estimate ITE as a whole and looks deeper into each of the $M$ subgroups to evaluate how each subgroup and the factors involved in the subgroup contribute to HTE estimation.
This modular, three-stage architecture of CRL is a deliberate design choice that prioritizes transparency, controllability, and clinical practicality, which are often more critical and fundamental than efficiency for generating trustworthy and actionable insights that can be integrated into clinical decision-making processes [kundu2021ai, shaban2020explainable]. Our workflow ensures that the entire process is interpretable and auditable, as each stage produces inspectable intermediate results. More importantly, it provides domain experts with multiple opportunities to intervene and guide the analysis, transforming the workflow from a black box into a collaborative, human-in-the-loop system. For instance, experts can pre-configure parameters (e.g., adjusting min.node.size to influence rule complexity) in the discovery phase, manually curate the initial candidate rule pool based on clinical knowledge after discovery, extend the ITE assumption to include non-linear interaction terms if theorized, and flexibly choose the depth of analysis applied to the selected rules.
Primarily due to the uncertainty and randomization of rule generation via causal forest, there may be multiple combinations of $\{\hat{\tau}_{m},\hat{r}_{m}\}^{M}_{m=1}$ generated in the rule discovery step, which may yield equivalent estimations of ITE. In this sense, our model is not identifiable. However, this problem is mitigated to a great extent by the subsequent selection step and analysis step. Besides, we can fit CRL repeatedly across resampled or perturbed datasets and identify rule combinations that consistently appear. This operation is recommended by [Bin2020Veridical, rewolinski2025pcs], which emphasizes the importance of reproducible and stable results under data and model perturbations. The results of our simulation studies in section 4 also verify its effectiveness. If, after such stability assessments, multiple rule combinations persist, this may reflect a genuine multiplicity in the underlying data-generating process and the inherent complexity of treatment effect heterogeneity. In such cases, the non-uniqueness of rule combinations is not a flaw of our method, but rather a strength. Our model thus serves as an exploratory tool for the complexity of HTE, and its final interpretation and validation should be guided by domain experts with prior knowledge and rigorous randomized trials.